OpenStudy (ellecullen):

HW help asap: A Cookie company makes three kinds of muffins; bran, chocolate chip, and blueberry. These are packaged in Regular, Extra, and Super Boxes. The regular box contains one dozen bran, one dozen chocolate chip. The Extra box contains two dozen bran, one dozen chocolate chip, and one dozen blueberry. The super box contains two dozen bran, two dozen chocolate chip, and the dozen blueberry. A total of 15dozen bran muffins, ten dozen chocolate chip muffins and eleven dozen blueberry muffins have been made for the day. Find the number of boxes of each that can be packaged without leftove

OpenStudy (ellecullen):

Sorry, it got cut off. without leftovers. For the purpose, define variables and write equations.

OpenStudy (lochana):

did you read it completely?

OpenStudy (lochana):

you can derive 3 equations for each muffin in terms of number of boxes

OpenStudy (lochana):

given that 15 bran, 10 chocolate , 11 blueberry require R(regular), E(extra) and S(super) number of boxes can be written as bran 15 = R + 2E + 2S chocolate 10 = R + E + 2S blueberry 11 = E + S

OpenStudy (lochana):

So E = 5 ; S = 6; R =-7

OpenStudy (lochana):

negative value of R means that you don't need regular boxes. I guess:)

OpenStudy (lochana):

yes. you don't need regular boxes.